Solve correctly step by step.
ID: A 23. A parallel plate capacitor consists of a movable metal plate and a fixed metal plate. The space between plates is filled with material whose dielectric constant is 4.0.The capacitor has a capacitance of2.0x 10 F b. If the charge on the capacitor is for the voltage across the capacitor. (8 points) 5 x 109 C, solve
Answer #1

Here ,

a)

let the plate area is A

2 *10^-10 = 4 * 8.854 *10^-12 * A/(1.5 *10^-4)

solving for A

A = 8.47 *10^-4 m^2

b)

voltage across the capacitor = (5 *10^-9)/(2 *10^-10)

voltage across the capacitor = 25 V